Given the population 1,2,3,5,7,11 and 13 compute the following.



A.population of mean



B.Population of variance



C.population standard deviation

1
Expert's answer
2022-06-06T08:19:27-0400

We have population values 1,2,3,5,7,11,13 population size N=7 .

A. Mean of population "(\mu)" = 


"\dfrac{1}{7}(1+2+3+5+7+11+13)=6"


B. Variance of population 


"\sigma^2=\dfrac{\Sigma(x_i-\bar{x})^2}{n}=\dfrac{1}{7}(25+16+9+1""+1+25+49)=\dfrac{126}{7}=18"


C.


"\sigma=\sqrt{\sigma^2}=\sqrt{18}=\sqrt{2}\approx4.24264"